# Decomposition theorems for unmatchable pairs in groups and field extensions

### v1

@@ Abstract.
A theory of matchings for finite subsets of an abelian group, introduced in connection with a conjecture of Wakeford on canonical forms for symmetric tensors, has since been extended to the setting of field extensions and to that of matroids. Earlier approaches have produced numerous criteria for matchability and unmatchability, but have offered little structural insight. In this paper, we develop parallel structure theorems which characterize unmatchable pairs in both abelian groups and field extensions. Our framework reveals analogous obstructions to matchability: nearly periodic decompositions of sets in the group setting correspond to decompositions of subspaces involving translates of simple extensions in the linear setting. This perspective not only recovers previously known results through short proofs, but also leads to new matching criteria and guarantees the existence of nontrivial unmatchable pairs.
1Department of Mathematics, Clayton State University, 2000 Clayton State Boulevard, Lake City, Georgia 92093, USA. maliabadi@clayton.edu.
@@
Keywords and phrases. Chowla sets and subspaces, Field extension, Matching property, Nearly periodic decomposition
@@ 1. Introduction
Overview. Let $G$ be an abelian group with operation written multiplicatively, and let $A,B\subseteq G$ be nonempty finite sets. A matching from $A$ to $B$ is a bijection $f:A\to B$ such that $af(a)\notin A$ for all $a\in A$. This notion was introduced by Fan and the second author of the present paper [9] in a work on Wakeford’s conjecture concerning canonical forms for symmetric tensors [14]. The approach taken in [9] involved the study of a specific class of matchings, called acyclic matchings, for certain pairs of subsets of $\mathbb{Z}^{n}$. When such matchings exist for all pairs $(A,B)$ satisfying obvious necessary conditions, the ambient group is said to have the acyclic matching property. This property was established for $\mathbb{Z}^{n}$ in [6], extended to torsion-free abelian groups in [12], and fully characterized for all abelian groups in [4].
Further progress has included broader results in abelian groups [1], an extension to arbitrary groups [7], enumerative results using the isoperimetric method [10], a linear formulation over field extensions [8], and a matroidal analogue [5]. For an application of matchings in the context of combinatorial number theory, see [11].
@@
Recently, in [3], the present authors established characterization theorems for matchable subsets in abelian groups and matchable subspaces over field extensions using Dyson transforms. These results removed a previous reliance on a diverse collection of number-theoretic inequalities and their linear analogues. The present paper continues that program by giving new structural characterizations of unmatchable sets and subspaces and exploring their implications, which include existence criteria for nontrivial unmatchable pairs.
@@
Organization of the paper. In Sections 1.1 and 1.2, we recall background information on matchings in abelian groups and field extensions. Section 2 establishes a structure theorem for unmatchable pairs $(A,B)$ of finite subsets of any abelian group (Theorem 2.3), exhibiting a nearly periodic decomposition for such pairs, and then develops consequences, e.g., the symmetric case $A=B$, a Chowla-type criterion, and inheritance of unmatchability to suitable quotients. Section 3 presents a linear analogue for field extensions: any unmatchable pair of subspaces decomposes as a direct sum built from translates of simple extensions (Theorem 3.3). Several corollaries follow, including results on self-matching and Chowla subspaces. Section 4 describes some directions for further work.
@@ 2. Structure theorem for unmatchable pairs in abelian groups
Below, we state and prove our structure theorem for unmatchable pairs $(A,B)$ in an abelian group. It will be used to derive new criteria for the existence and nonexistence of matchings, and it will serve as the basis for a new and concise proof of a known result concerning Chowla sets. An analogous result in the setting of field extensions will be presented in Section 3.
@@
Let $G$ be an abelian group with at least one finite nontrivial proper subgroup. Let $n$ be an integer such that $n_{0}(G)\leq n<|G|$ and $n\not\equiv n_{0}(G)-1\pmod{n_{0}(G)}$. Then there exist nonempty finite subsets $A,B\subseteq G$ such that $1\notin B$, $|A|=|B|=n$, and $(A,B)$ is an unmatchable pair.
@@
Let $G$ be an abelian group with at least one finite nontrivial proper subgroup. Let $n$ be an integer such that $n_{0}(G)\leq n<|G|$. Then there exist nonempty finite subsets $A,B\subseteq G$ such that $1\notin B$, $|A|=|B|=n$, and $(A,B)$ is an unmatchable pair if and only if there is a finite subgroup $H$ of $G$ with $|H|\leq n$ and $|H|\nmid(n+1)$.
@@ 3. Structure theorem for unmatchable subspaces in a field extension
The following theorem, a main result of [3], provides a framework for studying matchable subspaces in field extensions. Using it, one can recover all known matching criteria in this setting. We will apply it to obtain a structure theorem for unmatchable pairs $(A,B)$.
@@
The proposition below explains, in large part, the significance of the condition $\langle SR\rangle=S$ in Theorem 3.1. A proof is given in [3].
@@
Let $K\subseteq L$ be a field extension, let $n$ be a positive integer, let $S$ and $R$ be nonzero $K$-subspaces of $L$, and let $x\in R$. Assume that $\langle SR\rangle=S$ and $\dim S\leq n$. Then $x$ is algebraic over $K$ and in fact $K(x):K]\leq n$. Also, for each $a\in S$, we have $aK(x)\subseteq S$.
We next state and prove the principal result of this section, giving a structural decomposition for unmatchable pairs $(A,B)$ over a field extension $K\subsetneq L$. This theorem is a linear analog of Theorem 2.3. Here, translates of simple field extensions, i.e., $K$-subspaces of the form $aK(x)$ with $a,x\in L$, play a role similar to that of cosets in the group setting. We note that such subspaces were considered briefly in the matching context in [2].
@@
Let $K\subsetneq L$ be a field extension, and let $A$ and $B$ be two $n$-dimensional $K$-subspaces of $L$, with $n>0$. Then the pair $(A,B)$ is unmatchable if and only if there exists a nonzero subspace $R\subseteq B$ such that $A$ and $B$ can be written as
$A$ $=S\oplus Y,$
$B$ $=R\oplus Z,$
@@
$S=\sum_{\begin{subarray}{c}a\in S\end{subarray}}aK(x)$
for every $x\in R$, and $\dim Y<\dim R$.
Note: The conditions on the subspaces in the above decomposition ensure that $S$ is nonzero.
@@
By condition (b) and Proposition 3.2, we have $aK(x)\subseteq S$ for every $a\in S$ and $x\in R$, i.e., $S$ is a vector space over $K(x)$ for each $x\in R$. In other words,
$S=\sum_{a\in S}aK(x)$
for every $x\in R$. Note that $R\neq K$ by (c), and we may assume that $R\subseteq B$; otherwise, we can replace $R$ by $\pi_{B}(R)$, where $\pi_{B}$ is the projection map $B\oplus K\longrightarrow B$ along $K$, and conditions (a)–(c) will still hold. Thus, condition (c) gives us
@@
In the following table, we summarize the main information regarding unmatchable pairs in abelian groups and their linear analogues.
@@
Objects $A,B\subseteq G$, $|A|=|B|<\infty$ KK–subspaces $A,B\subseteq L$, $\dim A=\dim B<\infty$
@@
Structure of $S$ $S$ is a union of $\langle R\rangle$–cosets $S=\sum_{a\in S}aK(x)$ for all $x\in R$
@@
$S=\sum_{a\in S}aK(x)$
for all $x\in R$. It is easy to see that the sets $S,R,Y,Z$ satisfy all of the hypotheses of Theorem 3.3. Therefore AA is not matched to $B$. ∎
@@
In particular, if $B$ is a Chowla subspace, then $1\notin B$. The following result, originally proved in [3], is retrieved easily using Theorem 3.3.
@@
For all $0\neq a\in A$ and $0\neq x\in B$, we have
$\dim aK(x)=[K(x):K]>\dim B=\dim A,$
@@
Assume $K=\mathbb{Q}$ and $L=\mathbb{Q}(\sqrt[p]{2})$, where $p$ is a prime. Let $A$ be any $\mathbb{Q}$-subspace of $L$ of dimension p1p-1, and let $B$ be the trace-zero subspace of $\mathbb{Q}(\sqrt[p]{2})$ over $\mathbb{Q}$, which has dimension p1p-1. Then $B$ is a Chowla subspace. By Corollary 3.5, it follows that $A$ is matched to $B$.
The lemma below will help us to establish our next corollary. It will also be used in the proof of Theorem 3.12 later in this section.
Lemma 3.7.
Let $K\subseteq L$ be a field extension, and let $R$ and $S$ be $K$-subspaces of $L$ of finite dimension. Assume that $S$ is nonzero and $aK(x)\subseteq S$ for all $a\in S$ and $x\in R$. Then $aK(R)\subseteq S$ for all $a\in S$. Also, $K(R):K]<\infty$ and $\dim_{K}S$ is a positive multiple of $[K(R):K]$.
Proof.
Let $\{x_{1},\ldots,x_{t}\}$ be a $K$-basis for $R$, and let $i\in[t]$. By assumption, we can view $S$ as a vector space over $K(x_{i})$. Consider the equation
$\dim_{K}S=\dim_{K(x_{i})}S\cdot[K(x_{i}):K].$
Since $0<\dim_{K}S<\infty$, the above implies $K(x_{i}):K]<\infty$, and so $x_{i}$ is algebraic over $K$. We thus have $K(R)=K(x_{1},\ldots,x_{t})=K[x_{1},\ldots,x_{t}]$. It is now clear that $aK(R)\subseteq S$ for all $a\in S$, i.e., $S$ is a vector space over $K(R)$. In turn, this gives us
$\dim_{K}S=\dim_{K(R)}S\cdot[K(R):K],$
so that $K(R):K]<\infty$ and $\dim_{K}S$ is a positive multiple of $[K(R):K]$. ∎
@@
where $R$ is a nonzero subspace, $S=\sum_{a\in S}aK(x)$ for every $x\in R$, and $\dim Y<\dim R$. Hence, by Lemma 3.7,
$S=\sum_{a\in S}aK(R).$
Since we also have $1\notin A$, it follows that $K(R)\cap S=\{0\}$. Consider now the projection map $\pi:A\to Y$ along $S$, i.e., the $K$-linear map satisfying
$\ker\pi=S\qquad\text{and}\qquad\pi|_{Y}=\mathrm{id}_{Y}.$
Restrict $\pi$ to the subspace $T=K(R)\cap A$. Then
@@
$n_{0}(K,L)=\min\{[F:K]\mid K\subsetneq F\subseteq L,\;[F:K]<\infty\}.$
The following result, which can be viewed as a linear analogue of Corollary 2.7, guarantees the existence of nontrivial unmatchable pairs of sufficiently large dimension $n$, assuming that $n$ avoids a certain congruence class.
@@
Let $K\subsetneq L$ be a field extension with at least one proper, nontrivial intermediate field of finite $K$-dimension. Let $n$ be a positive integer such that $n_{0}(K,L)\leq n<[L:K]$ and $n\not\equiv n_{0}(K,L)-1\pmod{n_{0}(K,L)}$. Then there exist $K$-subspaces AA and $B$ of $L$ such that $1\notin B$, $\dim A=\dim B=n$, and the pair $(A,B)$ is unmatchable.
@@
Also, $\dim A=\dim B=n$, and $1\notin B$ since $1\notin R$ and $Z\cap K(x)=\{0\}$. Finally, for each $y\in R$, the field $K(y)$ is contained in $K(x)$, and so $S=\sum_{a\in S}aK(y)$. Therefore, by Theorem 3.3, the pair $(A,B)$ is unmatchable. ∎
@@
where for every $x\in R$,
$S=\sum_{a\in S}a\,\mathbb{Q}(x),$
and $\dim Y<\dim R$. Note that $S\neq\{0\}$. Also, we see that $\dim\mathbb{Q}(x)\leq\dim S\leq 3$ for all $x\in R$.
Now, the only nontrivial, proper intermediate subfield of $\mathbb{Q}\subseteq\mathbb{Q}(\sqrt[4]{2})$ is $\mathbb{Q}(\sqrt{2})$. Therefore, since $\{0\}\neq R\subseteq B$ and $1\notin B$, we have $\mathbb{Q}(x)=\mathbb{Q}(\sqrt{2})$ for some (in fact, all) nonzero $x\in R$. Hence
@@
On the other hand, since $\mathbb{Q}(x)=\mathbb{Q}(\sqrt{2})$ for every nonzero $x\in R$, it follows that $R\subseteq\mathbb{Q}(\sqrt{2})$, so $\dim R\leq 2$. Hence $\dim R=2$, and we now have $R=\mathbb{Q}(\sqrt{2})$, which contradicts the assumption that $1\notin B$.
Therefore, the pair $(A,B)$ is matchable.
@@
Let $K\subsetneq L$ be a field extension with at least one proper, nontrivial intermediate field of finite $K$-dimension. Let $n$ be a positive integer such that
@@
Assume there exist $n$-dimensional $K$-subspaces $A,B\subseteq L$ such that $1\notin B$ and $(A,B)$ is unmatchable. Applying Theorem 3.3, we obtain subspaces $R,S,Y,Z$ as described therein. Observe that since
$S=\sum_{a\in S}aK(x)$
for every $x\in R$, it follows that $S$ is a vector space over each field $K(x)$ for $x\in R$. Hence, by Lemma 3.7, $S$ is a vector space over $K(R)$ and $K(R):K]<\infty$.
Let $F=K(R)$ and $m=[F:K]$. Again by Lemma 3.7, the $K$-dimension of $S$ must be a positive multiple of $m$. We now show that $F$ can serve as the desired intermediate field of $K\subsetneq L$. Write $\dim S=mq$ where $q\geq 1$. Then
@@
Next we define $R$ and $S$. Choose a subspace $R\subseteq F$ with $\dim R=m-1$ and $R\cap K=\{0\}$. Note that $L:F]>n/m\geq q$, and let $a_{1},\dots,a_{q}\in L$ be linearly independent over $F$. Define

### v3

@@ Abstract.
A theory of matchings for finite subsets of an abelian group, introduced in connection with a conjecture of Wakeford on canonical forms for homogeneous polynomials, has since been extended to the setting of field extensions and to that of matroids. Earlier approaches have produced numerous criteria for matchability and unmatchability, but have offered little structural insight. In this paper, we develop parallel structure theorems which characterize unmatchable pairs in both abelian groups and field extensions. Our framework reveals analogous obstructions to matchability: nearly periodic decompositions of sets in the group setting correspond to decompositions of subspaces involving translates of a subfield in the linear setting. This perspective not only recovers previously known results through short proofs, but also leads to new matching criteria and guarantees the existence of nontrivial unmatchable pairs.
1Department of Mathematics, Clayton State University, 2000 Clayton State Boulevard, Lake City, Georgia 30260, USA. maliabadi@clayton.edu.
@@
Keywords and phrases. Chowla sets and subspaces, Field extension, Matching obstruction, Nearly periodic decomposition
@@ 1. Introduction
Overview. Let $G$ be an abelian group with operation written multiplicatively, and let $A,B\subseteq G$ be nonempty finite sets. A matching from $A$ to $B$ is a bijection $f:A\to B$ such that $af(a)\notin A$ for all $a\in A$. This notion was introduced in [9] in connection with a conjecture of Wakeford concerning the possible sets of monomials that can be removed from a generic homogeneous polynomial through a linear change in variables [13].
More specifically, to any set $A$ of same-degree monomials in the variables $x_{1},\ldots,x_{n}$ with |A|=n(n1)|A|=n(n-1), one associates a weighted bipartite graph $\mathcal{G}_{A}$. It was shown in [9] that the monomials in $A$ are generically removable if and only if the biadjacency matrix of $\mathcal{G}_{A}$ has nonzero determinant. The nonzero terms in the Leibniz expansion of this determinant correspond to certain matchings $A\longrightarrow B$, where the monomials in $A$ are identified with their exponent vectors and $B$ is a set of vectors corresponding to the operators $x_{i}\frac{\partial}{\partial x_{j}}$ for $i\neq j$. It turns out that when each monomial in $A$ involves all of the variables, there exists an “acyclic” matching corresponding to a determinant term that cannot be canceled. In this way, such matchings provide a combinatorial certificate for removability and hence a mechanism for studying canonical forms for homogeneous polynomials.
Acyclic matchings have been studied closely. When an acyclic matching exists for each pair $(A,B)$ of subsets of an abelian group $G$ with $|A|=|B|<\infty$ and $1\notin B$, one says that $G$ has the “acyclic matching property.” This property was established for $\mathbb{Z}^{n}$ in [6], extended to all torsion-free abelian groups in [11], and fully characterized for all abelian groups in [4]. Further progress has included broader results on matchings in abelian groups [1], an extension to arbitrary groups [7], a lower bound via Hamidoune’s isoperimetric method [10], a linear formulation over field extensions [8], and a matroidal analogue [5].
@@
Recently, in [3], the present authors established characterization theorems for matchable subsets in abelian groups and matchable subspaces over field extensions using Dyson transforms. These results removed a previous reliance on a diverse collection of number-theoretic inequalities and their linear analogues. We show here that the two main characterization theorems in [3] can be used to derive structural decomposition theorems for unmatchable pairs. Exploiting these decompositions, we obtain existence criteria for nontrivial unmatchable pairs and further consequences, both for finite subsets of abelian groups and for finite-dimensional subspaces over field extensions.
@@
Organization of the paper. In Sections 1.1 and 1.2, we recall background information on matchings in abelian groups and field extensions. Section 2 establishes a structure theorem for unmatchable pairs $(A,B)$ of finite subsets of any abelian group (Theorem 2.3), exhibiting a nearly periodic decomposition for such pairs, and then develops consequences, e.g., the symmetric case $A=B$, a Chowla-type criterion, and inheritance of unmatchability to suitable quotients. Section 3 presents a linear analogue for field extensions: any unmatchable pair of subspaces decomposes as a direct sum built in part from translates of a subfield (Theorem 3.3). Several corollaries follow, including results on self-matching and Chowla subspaces. Section 4 describes some directions for further work.
@@ 2. Structure theorem for unmatchable pairs in abelian groups
Below, we state and prove our structure theorem for unmatchable pairs. This result offers a different and complementary perspective to that of Theorem 2.1. It shows that unmatchability can be understood not just as the failure of a Hall-type condition, but as the consequence of a structural feature of the pair $(A,B)$. More precisely, unmatchability occurs when there exists a nonempty set $R\subseteq B$ such that $A$ is nearly a union of cosets of the subgroup $\langle R\rangle$, with a remainder set $Y$ of smaller size than $R$. This viewpoint yields concrete certificates of unmatchability, shows that several seemingly disparate matching results actually have a common underlying source, and serves as the main engine for new consequences in this paper (including existence criteria for nontrivial unmatchable pairs).
@@
We remark that the set $Z$ in a nearly periodic decomposition will play a passive role in this paper, receiving attention only when we are constructing unmatchable pairs $(A,B)$ with $1\notin B$, as we do in the proofs Corollary 2.7 and Theorem 2.9.
@@
Let $G$ be an abelian group with at least one finite nontrivial proper subgroup. Let $n$ be an integer such that $n_{0}(G)\leq n<|G|$ and $n\not\equiv n_{0}(G)-1\pmod{n_{0}(G)}$. Then there exist subsets $A,B\subseteq G$ such that $1\notin B$, $|A|=|B|=n$, and $(A,B)$ is an unmatchable pair.
@@
Let $G$ be an abelian group with at least one finite nontrivial proper subgroup. Let $n$ be an integer such that $n_{0}(G)\leq n<|G|$. Then there exist subsets $A,B\subseteq G$ such that $1\notin B$, $|A|=|B|=n$, and $(A,B)$ is an unmatchable pair if and only if there is a subgroup $H$ of $G$ with $|H|\leq n$ and $|H|\nmid(n+1)$.
@@ 3. Structure theorem for unmatchable subspaces in a field extension
Let $K\subseteq L$ be a field extension. Given a $K$-subspace $S$ of $L$, we will write $\dim_{K}S$ or $\dim S$ for its $K$-dimension, preferring the latter in situations where no confusion can arise.
The following theorem, a main result of [3], provides a framework for studying matchable subspaces in field extensions. We will apply it to obtain a structure theorem for unmatchable pairs $(A,B)$.
@@
The proposition below gives an indication of how the condition $\langle SR\rangle=S$ in Theorem 3.1 will be used in the rest of this section.
@@
Let $K\subseteq L$ be a field extension, let $n$ be a positive integer, and let $S$ and $R$ be nonzero $K$-subspaces of $L$. Assume that $\langle SR\rangle=S$ and $\dim_{K}S\leq n$. The following statements hold:
  • (i)
    Let $x\in R$. Then $aK(x)\subseteq S$ for all $a\in S$. In addition, $K(x):K]\leq n$, so that $x$ is algebraic over $K$.
  • (ii)
    We in fact have $aK(R)\subseteq S$ for all $a\in S$. Also, $K(R):K]\leq n$ and $\dim_{K}S$ is a positive multiple of $[K(R):K]$.
Proof.
Part (i) is proved in [3]. For (ii), observe that for any $a\in S\setminus\{0\}$, we have $\dim_{K}R=\dim_{K}aR$ and $aR\subseteq\langle SR\rangle=S$. Hence $\dim_{K}R\leq\dim_{K}S\leq n$. Let $\{x_{1},\ldots,x_{t}\}$ be a $K$-basis for $R$. According to (i), each basis element $x_{j}$ is algebraic over $K$, giving us $K(R)=K(x_{1},\ldots,x_{t})=K[x_{1},\ldots,x_{t}]$. We also have $aK(x_{j})\subseteq S$ for all $a\in S$ and $j\in[t]$, again by (i). It is now clear that $aK(R)\subseteq S$ for all $a\in S$. In other words, $S$ can be regarded as a vector space over $K(R)$.
For the rest of (ii), consider the equation
$\dim_{K}S=\dim_{K(R)}S\cdot[K(R):K].$
Since $0<\dim_{K}S\leq n<\infty$, it follows that $K(R):K]\leq n$ and $\dim_{K}S$ is a positive multiple of $[K(R):K]$. ∎
We are ready to state and prove one of the main results of this section, giving a structural decomposition for unmatchable pairs $(A,B)$ over a field extension $K\subsetneq L$. This theorem is a linear analogue of Theorem 2.3, but there is more to notice about it. In the field extension setting, matchability is controlled by the lattice of intermediate fields: unmatchability forces $A$ to be almost a sum of translates of an intermediate field $K(R)$ for some $K$-subspace $R$, with a remainder subspace $Y$ having smaller dimension than $R$. Equivalently, if we view $L$ as a vector space over $K(R)$ via multiplication, Theorem 3.3 shows that unmatchability forces a substantial portion of $A$ to be a $K(R)$-subspace of $L$, again with a small remainder subspace $Y$. This yields a structural characterization of the failure of matchability which depends on the extension $K\subsetneq L$ and leads to results that are most naturally formulated in terms of properties of intermediate extensions (for instance, Corollary 3.7 and Theorem 3.11).
@@
Let $K\subsetneq L$ be a field extension, and let $A$ and $B$ be two $n$-dimensional $K$-subspaces of $L$, with $n>0$. Then the pair $(A,B)$ is unmatchable if and only if there exists a nonzero $K$-subspace $R$ of $B$ such that $A$ and $B$ can be written as
$A=S\oplus Y,\qquad B=R\oplus Z,$
@@
$S=\sum_{\begin{subarray}{c}a\in S\end{subarray}}aK(R)$
and $\dim Y<\dim R$.
Note: Expressing $S$ in the above way is tantamount to saying that $S$ can be viewed as a vector space over $K(R)$. As such, $S$ has finite dimension, since $\dim_{K(R)}S\cdot[K(R):K]=\dim_{K}S\leq n$. Hence there exist $a_{1},\ldots,a_{m}\in S$ such that $S=\bigoplus_{i=1}^{m}a_{i}K(R)$. The dimension bound on $Y$ guarantees that $S\neq\{0\}$.
@@
By condition (b) and Proposition 3.2, we have $aK(R)\subseteq S$ for all $a\in S$. Thus
$S=\sum_{a\in S}aK(R).$
Note that $R\neq K$ by (c), and we may assume that $R\subseteq B$; otherwise, we can replace $R$ by $\pi_{B}(R)$, where $\pi_{B}$ is the projection map $B\oplus K\longrightarrow B$ along $K$, and conditions (a)–(c) will still hold. Hence condition (c) gives us
@@
In the following table, we summarize the main information regarding unmatchable pairs in abelian groups and over field extensions.
@@
Objects $A,B\subseteq G$, $0<|A|=|B|<\infty$ KK–subspaces $A,B\subseteq L$, $0<\dim A=\dim B<\infty$
@@
Structure of $S$ $S$ is a union of $\langle R\rangle$–cosets $S=\sum_{a\in S}aK(R)$
@@
$S=\sum_{a\in S}aK(R).$
It is easy to see that the sets $S,R,Y,Z$ satisfy all of the hypotheses of Theorem 3.3. Therefore AA is not matched to $B$. ∎
@@
In particular, if $B$ is a Chowla subspace, then $1\notin B$. The next result, originally proved in [3], is retrieved easily using Theorem 3.3.
@@
For any $a\in A\setminus\{0\}$, any nonzero $K$-subspace $R\subseteq B$, and any $x\in R\setminus\{0\}$, we have
$\dim aK(R)\geq\dim aK(x)=[K(x):K]>\dim B=\dim A,$
@@
Assume $K=\mathbb{Q}$ and $L=\mathbb{Q}(\sqrt[p]{2})$, where $p$ is a prime. Let $A$ be any $\mathbb{Q}$-subspace of $L$ of dimension p1p-1, and let $B$ be the trace-zero subspace of $\mathbb{Q}(\sqrt[p]{2})$ over $\mathbb{Q}$, which has dimension p1p-1. Then $B$ is a Chowla subspace. By Corollary 3.5, we conclude that $A$ is matched to $B$.
@@
where $R$ is a nonzero $K$-subspace, $S=\sum_{a\in S}aK(R)$, and $\dim Y<\dim R$. Observe that $K(R)\cap S=\{0\}$; otherwise, we can choose a nonzero $a\in K(R)\cap S$, and then $1=aa^{-1}\in aK(R)\subseteq S\subseteq A$, contradicting $1\notin A$. Consider now the projection map $\pi:A\to Y$ along $S$, i.e., the $K$-linear map satisfying $\ker\pi=S$ and $\pi|_{Y}=\mathrm{id}_{Y}$. Restrict $\pi$ to the subspace $T=K(R)\cap A$. Then
@@
$n_{0}(K,L)=\min_{F}\,[F:K],$
where $F$ ranges over all fields such that $K\subsetneq F\subseteq L$ and $F:K]<\infty$. The following result, which can be viewed as a linear analogue of Corollary 2.7, guarantees the existence of nontrivial unmatchable pairs of sufficiently large dimension $n$, assuming that $n$ avoids a certain congruence class.
@@
Let $K\subsetneq L$ be a field extension with at least one proper nontrivial intermediate field of finite $K$-dimension. Let $n$ be a positive integer such that $n_{0}(K,L)\leq n<[L:K]$ and $n\not\equiv n_{0}(K,L)-1\pmod{n_{0}(K,L)}$. Then there exist $K$-subspaces AA and $B$ of $L$ such that $1\notin B$, $\dim A=\dim B=n$, and the pair $(A,B)$ is unmatchable.
@@
Note that $1\notin B$, since $1\notin R$ and $Z\cap K(x)=\{0\}$. Also, $\dim A=\dim B=n$. Therefore, by Theorem 3.3, the pair $(A,B)$ is unmatchable. ∎
@@
where
$S=\sum_{a\in S}a\,\mathbb{Q}(R)$
and $\dim Y<\dim R$. Note that $S\neq\{0\}$. Also, we see that $\dim\mathbb{Q}(R)\leq\dim S\leq 3$.
Now, the only nontrivial proper intermediate subfield of $\mathbb{Q}\subseteq\mathbb{Q}(\sqrt[4]{2})$ is $\mathbb{Q}(\sqrt{2})$. Therefore, since $\{0\}\neq R\subseteq B$ and $1\notin B$, we have $\mathbb{Q}(R)=\mathbb{Q}(\sqrt{2})$. Hence
@@
On the other hand, since $\mathbb{Q}(R)=\mathbb{Q}(\sqrt{2})$, it follows that $\dim R\leq 2$. Hence $\dim R=2$, and we now have $R=\mathbb{Q}(\sqrt{2})$, which contradicts the assumption that $1\notin B$. Therefore, the pair $(A,B)$ is matchable.
@@
Let $K\subsetneq L$ be a field extension with at least one proper nontrivial intermediate field of finite $K$-dimension. Let $n$ be a positive integer such that
@@
Assume there exist $n$-dimensional $K$-subspaces $A,B\subseteq L$ such that $1\notin B$ and $(A,B)$ is unmatchable. Applying Theorem 3.3, we obtain subspaces $R,S,Y,Z$ as described therein, with
$S=\sum_{a\in S}aK(R).$
Note that $K(R):K]\leq\dim S\leq n$.
Let $F=K(R)$ and $m=[F:K]$. By Proposition 3.2, the $K$-dimension of $S$ must be a positive multiple of $m$. We now show that $F$ can serve as the desired intermediate field of $K\subsetneq L$. Write $\dim S=mq$ where $q\geq 1$. Then
@@
Next we define $R$ and $S$. Choose a $K$-subspace $R\subseteq F$ with $\dim R=m-1$ and $R\cap K=\{0\}$. Note that $L:F]>n/m\geq q$, and let $a_{1},\dots,a_{q}\in L$ be linearly independent over $F$. Define